if the measure of the angles of a triangle are in the ratio 2:3:4, the measure of the smallest angle is
1) 60 degrees
2) 20 degrees
3) 40 degrees
4) 80 degrees

Answer by user_dude2008(1862)

2x+3x+4x=180
9x=180
x=20
2x=2*20=40
Answer: 40